# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 3.1. From Transport to Variant Isoperimetry
$\theta_{\tau,\kappa}(\pi_{XW})$ $=\inf_{\pi_{Y|W=w}\in P_{\kappa},\forall w}\sup_{\lambda\geq 0}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W})$
(3.18) $\qquad+\lambda(\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})-\tau)$
$=\sup_{\lambda\geq 0}\inf_{\pi_{Y|W=w}\in P_{\kappa},\forall w}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W})$
(3.19) $\qquad+\lambda(\mathbb{E}_{\pi_{W}}[\mathsf{C}(\pi_{X|W}(\cdot|W),\pi_{Y|W}(\cdot|W))]-\tau)$
$=\sup_{\lambda\geq 0}\inf_{\pi_{Y|W=w}\in P_{\kappa},\forall w}\mathbb{E}_{\pi_{W}}[D(\pi_{Y|W}(\cdot|W)\|\nu_{Y})$
(3.20) $\qquad+\lambda(\sup_{f+g\leq c}\pi_{X|W}(f|W)+\pi_{Y|W}(g|W)-\tau)]$
$=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\inf_{\pi_{Y|W=w}\in P_{\kappa}}\sup_{f+g\leq c}D(\pi_{Y|W=w}\|\nu_{Y})$
(3.21) $\qquad+\lambda(\pi_{X|W=w}(f)+\pi_{Y|W=w}(g)-\tau)]$
$=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\sup_{f+g\leq c}\inf_{\pi_{Y|W=w}\in P_{\kappa}}D(\pi_{Y|W=w}\|\nu_{Y})$
(3.22) $\qquad+\lambda(\pi_{X|W=w}(f)+\pi_{Y|W=w}(g)-\tau)]$
$=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\sup_{f}\inf_{\gamma_{Y}\in P_{\kappa}}D(\gamma_{Y}\|\nu_{Y})$
(3.23) $\qquad+\lambda(\pi_{X|W=w}(f)+\gamma_{Y}(f^{c})-\tau)]$
where
  • the inf and sup are swapped in (3.19) which follows by the general minimax theorem [52, Theorem 5.2.2] together with the convexity of the relative entropy and the optimal transport cost functional;
  • (3.20) follows by the Kantorovich duality with $f,g$ denoting bounded continuous functions;
  • in (3.21) $\inf_{\pi_{Y|W}}$ is taken in a pointwise way;
  • the inf and sup are swapped in (3.22) which follows by the general minimax theorem [52, Theorem 5.2.2] by identifying that 1) the optimal value of the sup-inf in (3.22) is finite (since upper bounded by $\lambda(\mathsf{C}(\pi_{X|W=w},\nu_{Y})-\tau)$), and 2) by choosing $f,g$ as the zero function, the objective function turns to be $\pi_{Y|W=w}\mapsto D(\pi_{Y|W=w}\|\nu_{Y})-\lambda\tau$ whose sublevels are compact under the weak topology;
  • (3.23) follows since the optimal $g$ in (3.22) is
    $g(y)=f^{c}(y):=\inf_{x}c(x,y)-f(x).$
We consider the following optimization problem:
(3.24) $\inf_{\gamma_{Y}\in P_{\kappa}}D(\gamma_{Y}\|\nu_{Y})+\lambda\gamma_{Y}(g),$
where $g=f^{c}$ and $\lambda\geq 0$. Since adding $f$ by any constant will not change the objective function value in (3.23), without loss of generality, we assume $g(x_{0})=0$.
We now prove that the optimal value of the problem above remains the same for all $\kappa\geq\kappa_{0}$, where $\kappa_{0}$ is a constant independent of $\lambda,g$. To this end, we first bound $g$. By the triangle inequality, we have that
$g(y_{1})-g(y_{2})$ $\leq\sup_{x\in F}d^{p}(x,y_{1})-d^{p}(x,y_{2})\leq\left(d(x_{0},y_{1})+d(F)\right)^{p},$
and
$g(y_{1})-g(y_{2})$ $\geq\inf_{x\in F}d^{p}(x,y_{1})-d^{p}(x,y_{2})$
$\geq\left([d(x_{0},y_{1})-d(F)]^{+}\right)^{p}-\left(d(x_{0},y_{2})+d(F)\right)^{p},$
where $d(F)$ denotes the diameter of $F$ and $x]^{+}:=\max\{x,0\}$. Setting $y_{2}$ to the constant point $x_{0}$ yields that
$\left([d(x_{0},y)-d(F)]^{+}\right)^{p}-\left(d(F)\right)^{p}\leq g(y)-g(x_{0})$ $\leq\left(d(x_{0},y)+d(F)\right)^{p}.$
Since we assume $g(x_{0})=0$, we have
$g^{-}(y)\leq g(y)$ $\leq g^{+}(y),$
with
$g^{-}(y)$ $:=\left([d(x_{0},y)-d(F)]^{+}\right)^{p}-d(F)^{p},$
$g^{+}(y)$ $:=\left(d(x_{0},y)+d(F)\right)^{p}.$
The optimization problem in (3.23) is an information projection problem, whose solution is known to be $\gamma_{Y}^{(\lambda)}$ given by $\frac{{\text{\rm d}}\gamma_{Y}^{(\lambda)}}{{\text{\rm d}}\nu_{Y}}=\frac{e^{-\lambda g}}{\nu_{Y}(e^{-\lambda g})}$ [22]. Here for brevity, we use the notation $\nu(f)=\mathbb{E}_{\nu}[f]$ for a measure $\nu$ and a function $f$. For this distribution,
$\gamma_{Y}^{(\lambda)}(g^{q})=\frac{\nu_{Y}(g^{q}e^{-\lambda g})}{\nu_{Y}(e^{-\lambda g})}.$
Compute its derivative as
$\frac{{\text{\rm d}}}{{\text{\rm d}}\lambda}\gamma_{Y}^{(\lambda)}(g^{q})$ $=-\frac{\nu_{Y}(g^{q+1}e^{-\lambda g})\nu_{Y}(e^{-\lambda g})-\nu_{Y}(ge^{-\lambda g})\nu_{Y}(g^{q}e^{-\lambda g})}{\nu_{Y}(e^{-\lambda g})^{2}}$
$=-\left(\gamma_{Y}^{(\lambda)}(g^{q+1})-\gamma_{Y}^{(\lambda)}(g)\gamma_{Y}^{(\lambda)}(g^{q})\right)\leq 0,$
where the last inequality follows by the Harris inequality for the random variable $g(Y)$ with $Y\sim\gamma_{Y}^{(\lambda)}$ and increasing functions $t\mapsto t$ and $t\mapsto t^{q}$. Hence, $\gamma_{Y}^{(\lambda)}(g^{q})$ is decreasing in $\lambda\geq 0$. That is, $\gamma_{Y}^{(\lambda)}(g^{q})\leq\gamma_{Y}^{(0)}(g^{q})=\nu_{Y}(g^{q})\leq\nu_{Y}((g^{+})^{q})=:N.$ By the assumption $\nu_{Y}(d(x_{0},\cdot)^{q})<\infty$, it holds that $N<\infty$ since $g^{+}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$. So, $\gamma_{Y}^{(\lambda)}(g^{q})\leq N$ is bounded. Moreover, due to the fact that $g^{-}\leq g$, it holds that $\gamma_{Y}^{(\lambda)}((g^{-})^{q})\leq N$ is also bounded.
Since $g^{-}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$, we have $g^{-}(y)^{q}\geq(1-\epsilon)d(x_{0},y)^{q}-C$ for some positive constants $\epsilon,C$. Therefore,
$\gamma_{Y}^{(\lambda)}((g^{-})^{q})\geq(1-\epsilon)\gamma_{Y}^{(\lambda)}(d(x_{0},\cdot)^{q})-C,$
which implies that $\gamma_{Y}^{(\lambda)}(d(x_{0},\cdot)^{q})\leq\frac{N+C}{1-\epsilon}$ is bounded. Moreover, the upper bound $\frac{N+C}{1-\epsilon}$ is independent of $\pi_{XW},\lambda,f$. When $\kappa$ is set to any value greater than $\frac{N+C}{1-\epsilon}$, the optimal value $\theta(\tau,\kappa,\pi_{XW})$ does not depend on $\kappa$, i.e.,
$\theta(\tau,\kappa,\pi_{XW})$ $=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\sup_{f\in L^{1}(F)}\inf_{\gamma_{Y}}D(\gamma_{Y}\|\nu_{Y})$
$\qquad+\lambda(\pi_{X|W=w}(f)+\gamma_{Y}(f^{c})-\tau)]$
$=\theta(\tau,\infty,\pi_{XW}).$
The second part in the theorem follows since for the case $\nu=\nu_{X}=\nu_{Y}$, it is obvious that $\psi(\alpha,\tau)\leq\alpha$ for all $\alpha\geq 0,\tau\geq 0$, and hence $\mathrm{dom}\psi=\{\alpha\geq 0,\tau\geq 0\}$. Moreover, $\psi(\alpha,0)=\alpha$ for $\alpha\geq 0$, and $\psi(0,\tau)=0$ for $\tau\geq 0$, both of which coincide $\lim_{n\to\infty}E_{n}(\alpha,\tau)$. ∎
We next extend Theorem 9 to the case of non-compactly supported $\nu_{X}$.
Theorem 10 (Asymptotic Exponent for Metric-Valued Cost and Non-Compactly Supported $\nu_{X}$).
Assume that $\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $1\leq p<\infty$. Assume that
(3.25) $\mathbb{E}_{\nu_{X}}[e^{\lambda d^{p}(X,x_{0})}]$ $<\infty,\;\forall\lambda\in(0,\infty),$
$\mathbb{E}_{\nu_{Y}}[d^{q}(Y,x_{0})]$ $<\infty,$
for some $q>p$, and $x_{0}\in\mathcal{X}$. Then, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi$, (3.16) still holds. In particular, if further assume $\nu=\nu_{X}=\nu_{Y}$, then (3.16) holds for any $\alpha\geq 0,\tau\geq 0$.
Proof.
Our proof idea is as follows. By Sanov’s theorem, we may restrict most coordinates to compact sets without altering the exponential order of the set’s size, though the remaining coordinates cannot be constrained. Fortunately, the assumption in (3.25) ensures that the measure $\nu_{X}^{\otimes n}$ exhibits stronger concentration than Gaussian measures. This allows us to cover the majority of the mass of the remaining coordinates at only a negligible cost (i.e., in terms of distance). In other words, the isoperimetric exponent is determined by the one for compact sets which was already characterized in Theorem 9. We next present the detailed proof.
Since $\mathcal{X}=\mathcal{Y}$ is Polish, any probability measure on it is tight. So, for any $\epsilon\in(0,1)$, there is a compact set $F\subseteq\mathcal{X}$ such that $\nu_{X}(F^{c})\leq\epsilon$ and $\nu_{Y}(F^{c})\leq\epsilon$. Let $\mathbf{X}=(X_{1},...,X_{n})\sim\nu_{X}^{\otimes n}$ and $Z_{i}:=\mathbbm{1}_{B^{c}}(X_{i})$, $i\in[n]$. Then, $\mathbf{Z}\sim\mathrm{Bern}(\nu_{X}(F^{c}))^{\otimes n}$. By Sanov’s theorem, for any $\epsilon^{\prime}\in(\epsilon,1)$,
$\mathbb{P}\{\sum_{i=1}^{n}Z_{i}\geq n\epsilon^{\prime}\}\leq e^{-nD(\epsilon^{\prime}\|\nu_{X}(F^{c}))}\leq e^{-nD(\epsilon^{\prime}\|\epsilon)},$
where $D(\epsilon^{\prime}\|\epsilon)=\epsilon^{\prime}\ln\frac{\epsilon^{\prime}}{\epsilon}+(1-\epsilon^{\prime})\ln\frac{1-\epsilon^{\prime}}{1-\epsilon}$ is the binary relative entropy function, and the second inequality follows since $\epsilon\mapsto D(\epsilon^{\prime}\|\epsilon)$ is decreasing for $\epsilon<\epsilon^{\prime}$. Since $\epsilon\mapsto D(\epsilon^{\prime}\|\epsilon)$ goes to infinity as $\epsilon\downarrow 0$, given any $\epsilon^{\prime}>0$, we can choose $\epsilon$ small enough so that $D(\epsilon^{\prime}\|\epsilon)>\alpha$. For example, we can choose $\epsilon=\epsilon^{\prime}e^{-1/\epsilon^{\prime 2}}$ and choose $\epsilon^{\prime}$ small enough. For any measurable set $A$ such that $\nu_{X}^{\otimes n}(A)\geq e^{-n\alpha}$, it holds that
$\mathbb{P}\{\mathbf{X}\in A,\,\sum_{i=1}^{n}Z_{i}<n\epsilon^{\prime}\}$ $\geq\mathbb{P}\{\mathbf{X}\in A\}-\mathbb{P}\{\sum_{i=1}^{n}Z_{i}\geq n\epsilon^{\prime}\}\geq e^{-n\alpha}-e^{-nD(\epsilon^{\prime}\|\epsilon)}.$
@@
Given any $\delta>0$, for all sufficiently large $n$,
@@
(3.26) $\mathbb{P}\{\mathbf{X}\in A,\,\sum_{i=1}^{n}Z_{i}<n\epsilon^{\prime}\}\geq e^{-n(\alpha+\delta)}.$
For a subset $\mathcal{J}\subseteq[n]$, denote $C_{\mathcal{J}}$ as the event that $X_{i}\in F$ for $i\in\mathcal{J}$ and $X_{i}\in F^{c}$ for $i\in\mathcal{J}^{c}$. Then, (3.26) can be rewritten as
(3.27) $\mathbb{P}\{\mathbf{X}\in A\cap(\bigcup_{|\mathcal{J}|\geq n(1-\epsilon^{\prime})}C_{\mathcal{J}})\}\geq e^{-n(\alpha+o(1))}.$
On the other hand, there are ${n\choose\leq n\epsilon^{\prime}}:=\sum_{i=1}^{\left\lfloor n\epsilon^{\prime}\right\rfloor}{n\choose i}$ of sets $\mathcal{J}\subseteq[n]$ such that $|\mathcal{J}|\geq n(1-\epsilon^{\prime})$. Note that by Sanov’s theorem, ${n\choose\leq n\epsilon^{\prime}}\leq e^{nH(\epsilon^{\prime})}$, where $H(\epsilon^{\prime})=-\epsilon^{\prime}\ln\epsilon^{\prime}-(1-\epsilon^{\prime})\ln(1-\epsilon^{\prime})$ is the binary entropy function of $\epsilon^{\prime}$. Combining this with (3.27) yields that
(3.28) $\max_{|\mathcal{J}|\geq n(1-\epsilon^{\prime})}\mathbb{P}\{\mathbf{X}\in A\cap C_{\mathcal{J}}\}\geq e^{-n\alpha^{\prime}},$
where $\alpha^{\prime}:=\alpha+\delta+H(\epsilon^{\prime})$. Let $\mathcal{J}^{*}$ be the optimal $\mathcal{J}$ attaining the maximum in the above equation. Without loss of generality, we assume $\mathcal{J}^{*}=[n^{*}]$ for some $n^{*}\geq n^{\prime}:=\left\lceil n(1-\epsilon^{\prime})\right\rceil$. Let $n^{\prime\prime}=n-n^{\prime}$. We denote
(3.29) $\tilde{A}:=A\cap C_{\mathcal{J}^{*}}.$
In the following, we decompose a vector $\mathbf{x}=(\mathbf{u},\mathbf{v})\in\mathcal{X}^{n}$ into first $n^{\prime}$ components $\mathbf{u}\in\mathcal{X}^{n^{\prime}}$ and last $n^{\prime\prime}$ components $\mathbf{v}\in\mathcal{X}^{n^{\prime\prime}}$.
Denote for $\mathbf{v}_{0}=(x_{0},...,x_{0})\in\mathcal{X}^{n^{\prime\prime}}$,
(3.30) $B:=B_{r}(\mathbf{v}_{0})=\{\mathbf{v}\in\mathcal{X}^{n^{\prime\prime}}:d(\mathbf{v},\mathbf{v}_{0})\leq r\}.$
Then,

### v4

@@ 3.5.3. Step 3: Information-Theoretic Bound
$D(\mu_{\mathbf{Y}|J}\|\nu^{\otimes n_{1}}|\mu_{J})$ $=(1-\epsilon_{n})D(\hat{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\epsilon_{n}D(\tilde{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}})$
$\geq(1-\epsilon_{n})D(\hat{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}}).$
On the other hand,
$D(\mu_{\mathbf{Y}|J}\|\nu^{\otimes n_{1}}|\mu_{J})$ $=D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+D(\mu_{\mathbf{Y}|J}\|\mu_{\mathbf{Y}}|\mu_{J})$
$\leq D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\ln 2,$
where the inequality follows since $D(\mu_{\mathbf{Y}|J}\|\mu_{\mathbf{Y}}|\mu_{J})=I_{\mu}(J;\mathbf{Y})\leq H_{\mu}(J)\leq\ln 2.$ Hence,
(3.20) $D(\hat{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}})\leq\frac{D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\ln 2}{1-\epsilon_{n}}.$
By choosing $Q_{\mathbf{Y}|\mathbf{X}}$ in (3.18) as the feasible solution $\hat{Q}_{\mathbf{Y}|\mathbf{X}}$, we then have that
(3.21) $\log\nu^{\otimes n_{1}}(B^{\sqrt{n\tau_{1}}})\leq\inf\frac{D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\log 2}{1-\epsilon_{n}},$
where the infimum is taken over all transition probabilities $\mu_{Y|X}$ satisfying (3.19).
@@
Combining (3.16) and (3.21) yields that given $\delta>0$, for sufficiently large $n$ and sufficiently small $\epsilon$, it holds that
$-\log\nu^{\otimes n}(\tilde{A}_{0}^{\sqrt{n\tau_{0}}}\times B^{\sqrt{n\tau_{1}}})$
(3.22) $\leq\inf(1+\delta)\left(n_{0}D(\pi_{Y|W}\|\nu|\pi_{W})+D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})\right)+n\delta$
for some $(\pi_{XW},\mu_{\mathbf{X}})$ such that
$n_{0}D(\pi_{X|W}\|\nu_{F}|\pi_{W})+D(\mu_{\mathbf{X}}\|\nu^{\otimes n_{1}})\leq n\alpha^{\prime},\quad\alpha^{\prime}:=\alpha+2\delta,$
where the infimum is taken over all distributions $(\pi_{Y|XW},\mu_{Y|X})$ satisfying
$\mathbb{E}_{\pi}[d^{2}(X,Y)]$ $\leq\tau_{0}-\delta,$
$\mathbb{E}_{\pi}[\mathrm{Var}_{\pi}(d^{2}(X,Y)|X,W)]$ $\leq\kappa,$
and (3.19). Since $\tau_{0},\tau_{1}$ are arbitrary nonnegative numbers such that $\tau_{0}+\left(\sqrt{\tau_{1}}+2\sqrt{\delta}\right)^{2}\leq\tau$, optimizing the bound over all such $\tau_{0},\tau_{1}$ yields that the infimum can be taken over all distributions $(\pi_{Y|XW},\mu_{Y|X})$ satisfying
(3.23) $n_{0}\mathbb{E}_{\pi}[d^{2}(X,Y)]+\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},Y)|x_{i}]$ $\leq n\tau^{\prime}\;\mu_{\mathbf{X}}\textrm{-a.e.}$
(3.24) $\mathbb{E}_{\pi}[\mathrm{Var}_{\pi}(d^{2}(X,Y)|X,W)]$ $\leq\kappa,$
(3.25) $\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},Y)|x_{i})$ $\leq nM\;\mu_{\mathbf{X}}\textrm{-a.e.}$
where $\tau^{\prime}=\tau-5\delta-4\sqrt{\delta\tau}$.
3.5.4. Step 4: Heat Flow
We next construct a feasible solution to (3.22) by using heat flow. Let $\mu_{X_{t}|X=x}$ be the heat flow starting from the point $x$, which forms a transition probability. Let
$\pi_{WXX_{t}}=\pi_{WX}\mu_{X_{t}|X}$
and let $\pi_{X_{t}|W}$ the conditional probability induced by this joint probability. Let
(3.26) $\pi_{WXX_{t}}^{*}=\pi_{WX}\pi_{X_{t}|XW}^{*}$
where $\pi_{X_{t}|XW}^{*}$ is the optimal transport map attaining $\mathsf{W}_{2}(\pi_{X|W},\pi_{X_{t}|W}|\pi_{W})$. Let $\mu_{\mathbf{X}\mathbf{X}_{t}}=\mu_{\mathbf{X}}\mu_{X_{t}|X}^{\otimes n_{1}}$.
We choose $(\pi_{Y|XW},\mu_{Y|X})=(\pi_{X_{t}|XW}^{*},\mu_{X_{t}|X})$ and choose the time $t$ properly to satisfy (3.23). Let $\omega$ be a function given by
$\mathsf{W}_{2}^{2}(t):=n_{0}\mathsf{W}_{2}^{2}(\pi_{X|W},\pi_{X_{t}|W}|\pi_{W})+\mathsf{W}_{2}^{2}(\mu_{\mathbf{X}},\mu_{\mathbf{X}_{t}}),$
which is continuous in $t\geq 0$. By Remark 7, $\nu$ satisfies a log-Sobolev inequality with a constant $K>0$. Let $t_{0}=\frac{1}{2K}\ln\frac{\alpha^{\prime}}{\delta}$. Let $\tau^{\prime\prime}$ be a fixed number $\leq\tau$ which will be specified later. We choose the time $t=t_{0}$ if $\mathsf{W}_{2}(t_{0})\leq\sqrt{n\tau^{\prime\prime}}$, and choose $t$ as the smallest time $t_{1}$ such that $\mathsf{W}_{2}(t_{1})=\sqrt{n\tau^{\prime\prime}}$ if $\mathsf{W}_{2}(t_{0})>\sqrt{n\tau^{\prime\prime}}$. Hence, $\mathsf{W}_{2}(t)\leq\sqrt{n\tau^{\prime\prime}}$.
We next verify the two requirements—the expectation requirement and variance requirement in (3.23) and (3.24) for $(\mu_{X_{t}|X}\circ\mu_{X|W},\mu_{X_{t}|X})$, where $\mu_{X_{t}|X}\circ\mu_{X|W}$ denotes the resulting transition probability by cascading $\mu_{X|W}$ and $\mu_{X_{t}|X}$. We first focus on the expectation requirement. By the triangle inequality, we obtain for any $\mathbf{x},\mathbf{x}^{\prime}\in B$,
$\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\delta_{\mathbf{x}}\right)$ $\leq\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}}\right)+\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}},\delta_{\mathbf{x}}\right)$
(3.27) $\leq\sqrt{n\delta}+\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}},\delta_{\mathbf{x}}\right),$
where $\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}}\right)\leq\mathsf{W}_{2}\left(\delta_{\mathbf{x}},\delta_{\mathbf{x}^{\prime}}\right)\leq\sqrt{n\delta}$ follows by contraction of Wasserstein metric under $\mathrm{RCD}(0,\infty)$ given in [7, (6.1)] or [8, Corollary 3.18]. Taking expectation $\mathbb{E}_{\mathbf{X}^{\prime}\sim\mu_{\mathbf{X}}}$ and by Jensen’s inequality,
@@ 3.5.4. Step 4: Heat Flow
$\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\delta_{\mathbf{x}}\right)$ $\leq\mathbb{E}_{\mathbf{X}^{\prime}\sim\mu}[\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{X}^{\prime}},\delta_{\mathbf{x}}\right)]+\sqrt{n\delta}$